OpenStudy (anonymous):

4/5 subtracted by 1/5 in simpiest form

geerky42 (geerky42):

Because denominator is already same, you just subtract 1 from 4.

geerky42 (geerky42):

\(\dfrac{4}{5}-\dfrac{1}{5} = \dfrac{4-1}{5}=\ldots?\)

OpenStudy (anonymous):

so what is the anser im 9

OpenStudy (xguardians):

What is 4 - 1?

OpenStudy (anonymous):

3/5 i dont now i shurugle

geerky42 (geerky42):

Yes, 3/5 is correct answer.